# MMEdge: Accelerating On-device Multimodal Inference via Pipelined Sensing and Encoding

### v2

@@ 6.2. Experiment Setup
Devices. For evaluations on the UAV testbed, we implement MMEdge on NVIDIA Jetson Orin Nano (Corporation, 2024a). For evaluations on public datasets, we implement MMEdge on the NVIDIA Jetson Xavier NX (Corporation, 2024b), which features a 6-core ARM v8.2 64-bit CPU, a 384-core Volta GPU with 48 Tensor Cores, and 16 GB of shared LPDDR4x memory. This device is widely used in real-world edge AI applications, such as speech recognition and autonomous driving. Model training is conducted offline on a high-performance server equipped with dual AMD EPYC 7K62 processors (48 cores, 96 threads) and eight NVIDIA RTX 4090 GPUs, each with 24 GB of memory.
Implementation. For evaluations on the UAV testbed, we implement MMEdge in an end-to-end manner. For evaluations on public datasets, we follow prior works (Xu et al., 2020; Li et al., 2021) to simulate data collection and resource dynamics. Specifically, we run background multimodal data collection processes to emulate the sensing overhead on edge devices. During evaluations on public datasets, data are loaded while preserving the original frame rate by inserting fixed sleep intervals to simulate sensing delays. To simulate dynamic resource availability in real-world scenarios, we conduct experiments under various conditions by limiting CPU usage via Linux cgroup.
@@
  • Blocking Inference (Wang et al., 2023b), which waits for all modality inputs before making prediction, ensuring complete data at the cost of higher latency.
  • Non-Blocking Inference (Wu et al., 2024), which processes available modalities as they arrive, enabling early predictions to reduce latency at the cost of incomplete information.
  • Modality Gaiting (Hou et al., 2023), which dynamically selects input modalities before inference to reduce sensing and computation overhead. We implemented it by dynamically reducing the data rate of less informative modalities.
  • Imputation-Based Methods (Li et al., 2021; Wu et al., 2024), which imputes missing data of slower modalities to enhance inference robustness under asynchronous data inputs. We mplemente it by training a GAN to impute the features of the slow modality.
  • Model Selection (Rastikerdar et al., 2024), which dynamically chooses model branches based on input data complexity to balance accuracy and efficiency. We implemented it by dynamic model switch.
@@
Sensing and Model Configurations. Here we introduce the setting of different sensing and model configurations across datasets. For RGB, radar, and LiDAR modalities in our datasets, we employ three levels of model complexity using ResNet-18, ResNet-34, and ResNet-50. For audio, we design three CNN-based encoders with varying channel sizes and depths, named as small, medium and large. We use the facebook/OPT-125M (Zhang et al., 2022) model as the LLM for answer generation in the NuScenes-Mini-QA dataset. In our self-collected dataset, we use ResNet for radar and YOLO series models for RGB (Redmon et al., 2016). For sensing configurations, we vary the chunk size for audio (800, 1000, 1200). The RGB frame rates are set to 20, 25, and 29 FPS in LRW; 2, 6, and 12 FPS in NuScenes-Mini-QA; and 10, 20, and 30 FPS in our self-collected dataset. Radar sampling rates in our dataset are 5, 10, and 20 Hz, and LiDAR sampling rates in NuScenes-Mini-QA are 2, 10, and 20 Hz.
Figure 9. Detection results in different scenarios.
@@ 6.5.2. Performance under Varying Resource Availability
We also evaluate the robustness of MMEdge under different resource availability. To simulate different levels of computational availability on Nvidia edge devices, we cap CPU usage at 100% (high), 64% (medium), and 50% (low) using cpulimit (Angelo Marletta, [n. d.]). Figure 12(b) shows that MMEdge consistently achieves low latency across all CPU levels while maintaining high accuracy, which demonstrates its ability to adaptively select sensing and model configurations in response to dynamic resource availability.
@@ 6.6.1. Ablation study
Figure 13(a) shows the performance of MMEdge on the LRW dataset when disabling different design components, including the temporal aggregation module (T.A.), adaptive multimodal configuration (optim.), and cross-modal speculative skipping (Skip.). First, removing the temporal aggregation module results in a significant 19.5% drop in accuracy, indicating its effectiveness in maintaining accuracy performance for the pipelined sensing and encoding framework. When disabling speculative skipping module, the end-to-end latency increases from 137.67 ms to 192.03 ms, with the waiting time growing from 78.48 ms to 138.43 ms. This highlights that a significant portion of latency comes from waiting for slower modalities such as video. Similarly, disabling adaptive configuration also leads to increased latency, highlighting their roles in optimizing performance under dynamic conditions. On the LRW dataset, where the baseline accuracy is already high, the adaptive multimodal configuration primarily contributes to latency reduction rather than accuracy improvement, whereas in more diverse dataset (such as NuScenes-QA-Mini 11(b)) it shows clearer benefits by enhancing the latency–accuracy trade-off. These results collectively demonstrate that each component contributes to either accuracy or latency performance and is essential for achieving low-latency, high-accuracy inference.

### v5

@@ 6.2. Experiment Setup
Devices. For evaluations on the UAV testbed, we implement MMEdge on NVIDIA Jetson Orin Nano (jetson_orin). For evaluations on public datasets, we implement MMEdge on the NVIDIA Jetson Xavier NX (jetson_xavier), which features a 6-core ARM v8.2 64-bit CPU, a 384-core Volta GPU with 48 Tensor Cores, and 16 GB of shared LPDDR4x memory. This device is widely used in real-world edge AI applications, such as speech recognition and autonomous driving. Model training is conducted offline on a high-performance server equipped with dual AMD EPYC 7K62 processors (48 cores, 96 threads) and eight NVIDIA RTX 4090 GPUs, each with 24 GB of memory.
Implementation. For evaluations on the UAV testbed, we implement MMEdge in an end-to-end manner. For evaluations on public datasets, we follow prior works (xu2020approxdet; li2021low) to simulate data collection and resource dynamics. Specifically, we run background multimodal data collection processes to emulate the sensing overhead on edge devices. During evaluations on public datasets, data are loaded while preserving the original frame rate by inserting fixed sleep intervals to simulate sensing delays. To simulate dynamic resource availability in real-world scenarios, we conduct experiments under various conditions by limiting CPU usage via Linux cgroup.
@@
  • Blocking Inference (wang2023patch), which waits for all modality inputs before making prediction, ensuring complete data at the cost of higher latency.
  • Non-Blocking Inference (wu2024adaflow), which processes available modalities as they arrive, enabling early predictions to reduce latency at the cost of incomplete information.
  • Modality Gaiting (hou2023smg), which dynamically selects input modalities before inference to reduce sensing and computation overhead. We implemented it by dynamically reducing the data rate of less informative modalities.
  • Imputation-Based Methods (li2021low; wu2024adaflow), which imputes missing data of slower modalities to enhance inference robustness under asynchronous data inputs. We mplemente it by training a GAN to impute the features of the slow modality.
  • Model Selection (rastikerdar2024cactus), which dynamically chooses model branches based on input data complexity to balance accuracy and efficiency. We implemented it by dynamic model switch.
@@
Sensing and Model Configurations. Here we introduce the setting of different sensing and model configurations across datasets. For RGB, radar, and LiDAR modalities in our datasets, we employ three levels of model complexity using ResNet-18, ResNet-34, and ResNet-50. For audio, we design three CNN-based encoders with varying channel sizes and depths, named as small, medium and large. We use the facebook/OPT-125M (zhang2022opt) model as the LLM for answer generation in the NuScenes-Mini-QA dataset. In our self-collected dataset, we use ResNet for radar and YOLO series models for RGB (redmon2016you). For sensing configurations, we vary the chunk size for audio (800, 1000, 1200). The RGB frame rates are set to 20, 25, and 29 FPS in LRW; 2, 6, and 12 FPS in NuScenes-Mini-QA; and 10, 20, and 30 FPS in our self-collected dataset. Radar sampling rates in our dataset are 5, 10, and 20 Hz, and LiDAR sampling rates in NuScenes-Mini-QA are 2, 10, and 20 Hz.
@@ 6.5.2. Performance under Varying Resource Availability
We also evaluate the robustness of MMEdge under different resource availability. To simulate different levels of computational availability on Nvidia edge devices, we cap CPU usage at 100% (high), 64% (medium), and 50% (low) using cpulimit (cpulimit). Figure 12(b) shows that MMEdge consistently achieves low latency across all CPU levels while maintaining high accuracy, which demonstrates its ability to adaptively select sensing and model configurations in response to dynamic resource availability.
@@ 6.6.1. Ablation study
Figure 13(a) shows the performance of MMEdge on the LRW dataset when disabling different design components, including the temporal aggregation module (T.A.), adaptive multimodal configuration (optim.), and cross-modal speculative skipping (Skip.). First, removing the temporal aggregation module results in a significant 19.5% drop in accuracy, indicating its effectiveness in maintaining accuracy performance for the pipelined sensing and encoding framework. When disabling speculative skipping module, the end-to-end latency increases from 137.67 ms to 192.03 ms, with the waiting time growing from 78.48 ms to 138.43 ms. This highlights that a significant portion of latency comes from waiting for slower modalities such as video. Similarly, disabling adaptive configuration also leads to increased latency, highlighting their roles in optimizing performance under dynamic conditions. On the LRW dataset, where the baseline accuracy is already high, the adaptive multimodal configuration primarily contributes to latency reduction rather than accuracy improvement, whereas in more diverse dataset (such as NuScenes-QA-Mini 9(b)) it shows clearer benefits by enhancing the latency–accuracy trade-off. These results collectively demonstrate that each component contributes to either accuracy or latency performance and is essential for achieving low-latency, high-accuracy inference.